OpenStudy (anonymous):

Find the solution of this system equations. x-7y=-20 and -3x+9y-36 Helpp?

OpenStudy (anonymous):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

I have no i idea.. The question only says what it says above.. I really don't understand Math at all. :/

OpenStudy (anonymous):

lets try substitution because it is a little easier. x - 7y = -20 --> x = 7y - 20 now sub 7y - 20 in for x in the other equation -3x + 9y = -36 -3(7y - 20) + 9y = -36 -- distribute through the parenthesis -21y + 60 + 9y = -36 -- subtract 60 from both sides -21y + 9y = -36 - 60 -- combine like terms - 11y = -96 -- divide by -11 y = -96/-11 y = 96/11 now sub 96/11 in for y in either of the equations x - 7y = -20 x - 7(96/11) = -20 x - 672/11 = -20 x = -20 + 672/11 x = - 220/11 + 672/11 x = 452/11 check... x - 7y = -20 452/11 - 7(96/11) = -20 452/11 - 672/11 = -20 -220/11 = -20 -20 = -20 (correct) x = 452/11 and y = 96/11 sorry it took so long, I had a little trouble

OpenStudy (anonymous):

It's fine. Thank you! (: